# Differentiable Optimization of Similarity Scores Between Models and Brains

### v1

@@ Abstract
What metrics should guide the development of more realistic models of the brain? One proposal is to quantify the similarity between models and brains using methods such as linear regression, Centered Kernel Alignment (CKA), and angular Procrustes distance. To better understand the limitations of these similarity measures we analyze neural activity recorded in five experiments on nonhuman primates, and optimize synthetic datasets to become more similar to these neural recordings. How similar can these synthetic datasets be to neural activity while failing to encode task relevant variables? We find that some measures like linear regression and CKA, differ from angular Procrustes, and yield high similarity scores even when task relevant variables cannot be linearly decoded from the synthetic datasets. Synthetic datasets optimized to maximize similarity scores initially learn the first principal component of the target dataset, but angular Procrustes captures higher variance dimensions much earlier than methods like linear regression and CKA. We show in both theory and simulations how these scores change when different principal components are perturbed. And finally, we jointly optimize multiple similarity scores to find their allowed ranges, and show that a high angular Procrustes similarity, for example, implies a high CKA score, but not the converse.
@@ 1 Introduction
Natural intelligence entails the interactions between many systems - such as different sensory, memory, and motor systems. Therefore, ultimately, many of the open questions in systems neuroscience will require models that bridge these systems. However, there are many challenges as we work towards these multi-system models of the brain. We need guidance to search through this huge design space of potential models. It would be helpful to have metrics that allow us to extract general principles that are applicable across many models, or alternatively, highlight the power of heterogeneity, and shine a spotlight on combinations of tasks, datasets, and models that are not currently well explained.
Similarity measures have become a cornerstone in evaluating representational alignment across different models (Kornblith et al., 2019), different biological systems (Kriegeskorte et al., 2008), and across both artificial and biological systems. Researchers have employed diverse methods to compare model representations with, for example, brain activity, aiming to identify models that exhibit brain-like representations (Yamins et al., 2014; Sussillo et al., 2015; Schrimpf et al., 2018; Nayebi et al., 2018). However, while these measures are actively used and provide an efficient way to compare structure across complex systems, it is not clear that they adequately represent the computational properties of interest, and there is a need to better understand their limitations. The field lacks clear guidelines for selecting the most appropriate measure for a given context.
In this work we study several popular methods that have been proposed to quantify the similarity between models and neural data, in particular, linear regression (Yamins et al., 2014; Schrimpf et al., 2018), Centered Kernel Alignment (CKA) (Kornblith et al., 2019), and angular Procrustes distance (Williams et al., 2021; Ding et al., 2021). See appendix B.1 for a brief overview. We analyzed neural data from studies on nonhuman primates (Figure 1) and compared the neural responses to task-optimized recurrent neural networks (RNNs), or synthetic datasets, with different similarity scores. In order to study what drives high similarity scores we directly optimize the synthetic datasets to maximize their similarity to the neural datasets as assessed by different methods, for example, linear regression, CKA, or angular Procrustes distance.
Comparing similarity scores across studies is challenging, primarily due to variability in naming and implementation conventions. As part of our contribution to the research community we have created, and are continuing to develop, a Python package that benchmarks and standardizes similarity measures.
@@
Figure 2: Different similarity measures are not consistent when comparing task-optimized recurrent neural networks to neural datasets. We consider two neural datasets from prefrontal cortex (PFC) (Mante et al., 2013) (a) and Frontal Eye Field (FEF) (Siegel et al., 2015) (b) in monkeys performing an experimental task that required the animal to attend to either color or motion information while ignoring the non-cued feature of the stimuli. On each trial, a field of colored moving dots is shown. Monkeys are given a cue at the beginning of the trial to determine whether the dots in the stimulus are moving left vs right, or are red vs green. The monkey reported its choice with a saccade to one of two visual targets. In both datasets, we analyzed neural activity taken when the dot stimulus was presented. In panel a neural activity is visualized in a low-dimensional space capturing task-relevant dynamics using the targeted dimensionality reduction method from Mante et al. 2013. Each curve shows the average neural activity for a different experimental condition. See Mante et al. 2013 for a detailed description of the analysis. This visualization highlights features of the data but the similarity scores were computed using the firing rates from the electrode recordings before any dimensionality reduction. In panel b neural firing rates for two example neurons are shown with the colors denoting average activity for different experimental conditions. (c, d) RNNs with three different architectures, CTRNN, LowPassCTRNN, LSTM and three different nonlinearities, ReLU, ReTanh, Tanh are compared to neural datasets. The Neuron-split baseline score was obtained by dividing the neurons from a single dataset into disjoint sets and then comparing. If the model-data scores are equal to the neuron-split scores this indicates that model activity is indistinguishable from the neural activity of other recorded neurons. The Condition-average baseline score was obtained by averaging the neural activity along the trial and condition dimensions and comparing it to the original dataset. Each neuron in this condition-averaged dataset still has a unique time-varying firing rate. This strong baseline shows the similarity to the original data that one can obtain by only keeping the condition-independent neural dynamics. The different similarity measures do not agree on the relative rankings of the models, or on a more global scale, how similar the models are relative to the two baseline scores.
@@ 2 Related Work
Reviews. Recent reviews by Sucholutsky et al. (2023) and Klabunde et al. (2023) provide comprehensive overviews of representational similarity measures and their theoretical properties. While these reviews highlight the diversity of available metrics, they offer limited practical guidance on metric selection.
Our work addresses this gap by proposing a general framework for comparing similarity measures on experimental neural datasets. By directly optimizing synthetic datasets to maximize their similarity to neural recordings, we can systematically investigate how different metrics prioritize various aspects of the data, such as specific principal components or task-relevant information.
@@
However, our approach differs significantly in how we construct a diverse set of datasets with varying behavioral characteristics. Instead of relying on pre-trained models, we begin with unstructured noise and directly optimize it to maximize similarity to neural recordings. This allows us to address questions such as whether high similarity scores can be achieved without encoding task-relevant variables. Our optimization-based approach reveals how different similarity measures guide the emergence of task-relevant information within the synthetic datasets, offering a dynamic perspective on their properties.
@@ 3 Method
To evaluate the similarity of representations between two systems, we extract feature representations such as activity in a brain area or model layer in response to stimuli. Our objective is to quantify the alignment between these representations using a similarity score. Assume two datasets $X$ and $Y$ represent these features with dimensions (sample, feature) and mean-centered columns. Datasets with temporal dynamics are reshaped from (time, sample, feature) to (time*sample, feature). We define a scoring function $\text{score}(X,Y)$ as a measure that increases with similarity, achieving a maximum of 1 when $X=Y$.
We show results for the following three common similarity measures (see appendix B.1 for details):
  • Linear Regression finds the best linear mapping $B$ that predicts a reference dataset $X$ from a dataset $Y$ (Yamins et al., 2014; Schrimpf et al., 2018). We measure the goodness of fit using $R^{2}$ and use ridge regularization with parameter $\lambda=100$ as well as 5-fold cross-validation that tests generalization across different experimental conditions.
    $R^{2}_{LR}=1-\dfrac{\min_{B}\|X-YB\|^{2}_{F}+\lambda\|B\|_{F}}{\|X\|^{2}_{F}}$
@@
    We also consider the angular version of CKA (Williams et al., 2021; Lange et al., 2022).
@@
    $\max_{Q\in\mathcal{O}}\dfrac{\langle X,QY\rangle}{\|X\|_{F}\|Y\|_{F}}$
@@ 4 Results
4.1 High similarity scores do not guarantee encoding of task relevant variables
Question. We start by asking if synthetic datasets with high similarity scores relative to the neural data, encode task relevant variables, for example, the stimulus features or the response of the monkey, in the same way as the neural data. More specifically, is it possible for the synthetic datasets to have a high similarity score while failing to encode task relevant variables?
Problem with CKA and linear regression. Surprisingly we find that for linear regression and CKA the answer is yes, a high similarity score does not necessarily mean the synthetic datasets encode task relevant variables like the neural data. Figures 3a and 3b show the decode accuracy of a linear classifier trained to decode task relevant variables (cross-validated across different conditions) as the similarity score increases. Before optimization, the synthetic datasets initially consisted of Gaussian noise and the decode accuracy was near the baseline chance level of 0.5 as expected for the binary classifier used in this analysis.
Consider the synthetic data optimized towards the Siegel 2015 neural recordings using CKA similarity (second row and column in Figure 3). When the synthetic dataset has a high similarity score of 0.9 the decode accuracy for all the task variables is still less than that found in the neural activity (horizontal dashed lines). This is in contrast to the case where synthetic data is optimized to maximize angular Procrustes similarity (first column) and a similarity score of near 0.9 yields a dataset that encodes task variables to the same degree as the neural recordings. Note that both CKA and angular Procrustes have the same similarity scale ranging between 0 and 1 (perfect similarity). See appendix B.1 for more details.
As another example, consider the decode accuracy for the color of the stimulus that was shown in the experiments (green curves). For linear regression and CKA the similarity score can be high even when the data does not encode this task variable at the level found in the neural activity, suggesting there is a mismatch between the datasets with a high similarity score and features of the neural activity. In contrast, for the angular Procrustes distance the synthetic datasets that have a high similarity score relative to the neural activity also encode the task relevant variables.
How high should a good score be? As a final more optimistic note, for all the similarity measures, we can see a general trend that optimizing the synthetic data towards higher similarity eventually leads the data to encode task relevant variables above the baseline chance level of 0.5 accuracy. One perspective on this decoding analysis is that it gives a window into what is a high similarity score. If a high score is one where the data is encoding task relevant variables then these similarity numbers will need to be interpreted differently based on the similarity measure.
High variability in the condition average baseline across metrics. A specific number for the similarity score, 0.9 for example, can have different interpretations depending on the measure that is used, as we have seen in the examples above contrasting CKA and angular Procrustes. Conversely if two models or datasets are compared then this comparison may result in different scores depending on the similarity measure. Prior work has used the condition average baseline as a reference score for benchmarking models against neural data (Cloos et al., 2022). The condition average baseline is obtained by averaging each neuron’s activity across trials from all experimental conditions, while still allowing each neuron to have its own unique time-varying firing rate, and then comparing this condition-averaged dataset to the original neural data. This baseline shows the similarity one can obtain by discarding the condition specific activity to keep only the condition independent dynamics of the neural activity. The condition average baseline is shown in Figures 3a and 3b as vertical dashed lines. For a single dataset, we are always comparing the same two quantities to obtain the condition average baseline but the scores are quite different depending on the similarity measure. For the angular Procrustes measure, as opposed to the linear regression similarity measure, this baseline score is closer to the point where task relevant information starts to be encoded.
Figure 3: (a, b) Decode accuracy for experimental variables versus similarity scores. Decode is from synthetic data optimized towards greater similarity with the neural data from (a) Mante et al. (2013) and (b) Siegel et al. (2015). Horizontal dashed lines indicate the decode accuracy from the neural data. (c, d) Decode accuracy from neural data versus number of principal components included in the decode. Decode uses data from (c) prefrontal cortex from Mante et al. 2013 and (d) FEF from Siegel et al. 2015.
@@ 4.2.1 Low-dimensional synthetic datasets
Question. How much of the neural data must be captured by a synthetic dataset or model before the decode accuracy reaches the level seen in the neural dataset itself? One perspective on this question is to decode the task variables from neural data after projecting onto principal components 1 through N, where principal component 1 captures the most variance. In order to capture all the information about the task variables, at least several principal components must be included in the decode as shown in Figures 3c and 3d. This motivates the following hypothesis.
@@
Hypothesis. Perhaps the reason that linear regression and CKA similarity scores can be so high while the synthetic data fails to encode task variables is because these similarity measures preferentially rely on the top few principal components.
We explore this hypothesis in the following set of analyses with a synthetic dataset based on the neural recordings from Mante et al. 2013.
Figure 4a shows the reference dataset (compare to Figure 2a). We can think of this reference dataset as a low-dimensional neural trajectory summarizing the population activity of many neurons, or alternatively, as the firing rates of two neurons over time (shown here encoding the two task variables of choice and dot motion coherence), recorded during six different experimental conditions, with the color in Figure 4a denoting the condition.
Figure 4b shows the transformation of an initially random Gaussian noise dataset as it is optimized to maximize either the angular Procrustes or CKA similarity score with respect to the reference dataset. The score increases from an initial value near 0 to a maximum near 1 as optimization progresses, with the insets at the top of the figure showing the optimized noise dataset at various points during this procedure. The yellow curve shows how well the optimized dataset captures the first principal component of the reference dataset, as quantified by $R^{2}$, throughout optimization (see appendix B.2 for details). Notice that the second principal component, shown in purple, is only captured at a much higher optimization score for CKA versus angular Procrustes.
Figure 4c shows the same results when a synthetic Gaussian noise dataset is optimized towards the reference dataset using either linear regression similarity or angular CKA similarity (Williams et al., 2021).
@@
Figure 4: Different similarity measures differentially prioritize learning principal components of the data. (a) Reference dataset used as a target during optimization. (b, c) Initial Gaussian random noise data is updated to maximize similarity with the reference dataset, as quantified by one of the similarity measures. The transformation of the random noise dataset is shown at the top of panel b. The first principal component of the reference dataset is increasingly well captured by the optimized data as the similarity scores increase (yellow curves). The second, lower variance, component is also learned when maximizing the angular Procrustes similarity but is only captured at high similarity scores when maximizing linear regression, CKA, and angular CKA similarity. (d) Four reference datasets with decreasing variance along the second principal component. (e) Similarity measures capture both principal components when their variance is approximately equal. However, when the variance differs, CKA and linear regression have a much more pronounced preference for learning the high variance component as opposed to angular Procrustes which causes the optimized dataset to also capture the low variance dimension (curves colored according to asymmetry of variance distribution).
@@
Dependence on the variance distribution. The optimization dynamics not only depend on the similarity measure but also on the variance distribution of the reference dataset. Figure 4d shows four reference datasets with the same variance along the first principal component but decreasing variance along the second. If both dimensions have approximately equal variance then angular Procrustes, CKA, and linear regression will learn both dimensions similarly during optimization as shown by the white curve in Figure 4e. The curves are colored to indicate the fraction of variance the second principal component has relative to the first, so 1 indicates both principal components have the same variance. As the asymmetry between the principal components grows, the optimization to maximize angular Procrustes similarity still prioritizes the lower variance principal component (first column, second row), while CKA and linear regression do not capture the second principal component of the reference dataset until much later during optimization.

### v3

@@ Abstract
How do we know if two systems – biological or artificial – process information in a similar way? Similarity measures such as linear regression, Centered Kernel Alignment (CKA), Normalized Bures Similarity (NBS), and angular Procrustes distance, are often used to quantify this similarity. However, it is currently unclear what drives high similarity scores and even what constitutes a “good” score. Here, we introduce a novel tool to investigate these questions by differentiating through similarity measures to directly maximize the score. Surprisingly, we find that high similarity scores do not guarantee encoding task-relevant information in a manner consistent with neural data; and this is particularly acute for CKA and even some variations of cross-validated and regularized linear regression. We find no consistent threshold for a good similarity score – it depends on both the measure and the dataset. In addition, synthetic datasets optimized to maximize similarity scores initially learn the highest variance principal component of the target dataset, but some methods like angular Procrustes capture lower variance dimensions much earlier than methods like CKA. To shed light on this, we mathematically derive the sensitivity of CKA, angular Procrustes, and NBS to the variance of principal component dimensions, and explain the emphasis CKA places on high variance components. Finally, by jointly optimizing multiple similarity measures, we characterize their allowable ranges and reveal that some similarity measures are more constraining than others. While current measures offer a seemingly straightforward way to quantify the similarity between neural systems, our work underscores the need for careful interpretation. We hope the tools we developed will be used by practitioners to better understand current and future similarity measures.
@@ 1 Introduction
Similarity measures have become a cornerstone in evaluating representational alignment across different models (Kornblith et al., 2019), different biological systems (Kriegeskorte et al., 2008b), and across both artificial and biological systems. Researchers have employed diverse methods to compare model representations with, for example, brain activity, aiming to identify models that exhibit brain-like representations (Yamins et al., 2014; Sussillo et al., 2015; Schrimpf et al., 2018; Nayebi et al., 2018). However, while these measures are actively used and provide an efficient way to compare structure across complex systems, it is not clear that they adequately represent the computational properties of interest, and there is a need to better understand their limitations. Whenever we choose a similarity measure, we are making a commitment to what we care about in the two systems we are comparing. Therefore, understanding what drives these similarity scores is crucial for understanding this commitment. The field lacks clear guidelines for interpreting similarity scores in a given experimental context.
In this work we study several popular methods that have been proposed to quantify the similarity between models and neural data, in particular, linear regression (Yamins et al., 2014; Schrimpf et al., 2018), Centered Kernel Alignment (CKA) (Kornblith et al., 2019), angular Procrustes distance (Williams et al., 2021; Ding et al., 2021), and Normalized Bures Similarity (NBS) (Tang et al., 2020). See appendix C.1 for a brief overview. We analyzed neural data from studies on nonhuman primates (Figure 1) and compared the neural responses to task-optimized recurrent neural networks (RNNs), or synthetic datasets, with different similarity scores. In order to study what drives high similarity scores we directly optimize the synthetic datasets to maximize their similarity to the neural datasets as assessed by different similarity measures.
@@
Our main contributions and findings can be summarized as follows:
  1. 1.
    We show that a good value for a similarity score varies depending on the similarity measure and the dataset (Figure 3). Furthermore, we demonstrate that a high similarity score does not guarantee that models encode task-relevant information in a manner consistent with neural data. This is particularly relevant for the many studies that rely on these similarity measures to quantitatively characterize the degree of alignment between models and brains.
  2. 2.
    We identify what drives high similarity scores by differentiating through the similarity measures to directly maximize the score. We discover that different similarity measures differentially prioritize learning principal components of the data. For example, CKA, as opposed to angular Procrustes, may indicate a high score when many of the lower variance components are not captured, even when these dimensions carry crucial task-related information (Figures 4 and 5).
  3. 3.
    Through theoretical derivation we show the sensitivity of CKA, angular Procrustes, and Normalized Bures Similarity to the variance of principal component dimensions, and explain the dependence CKA shows to high variance components (section 4.3 and Figure 6)
  4. 4.
    We characterize the allowable range of scores between two different similarity measures by jointly optimizing their scores. Surprisingly, we reveal that a high angular Procrustes similarity implies a high CKA score, but not the converse (for the dataset shown in Figure 7).
  5. 5.
    Comparing similarity scores across studies is challenging, primarily due to variability in naming and implementation conventions. As part of our contribution to the research community we have created, and are continuing to develop, a Python package that benchmarks and standardizes similarity measures. Currently there are approximately 100 different similarity measures from 14 packages. Similarity package: https://github.com/nacloos/similarity-repository
Figure 2: Different similarity measures do not agree on the relative rankings when comparing models to neural datasets. One example application of similarity measures is to evaluate the similarity of task-optimized recurrent neural networks to neural datasets. We consider two neural datasets from (a) prefrontal cortex (PFC) (Mante et al., 2013) and (b) Frontal Eye Field (FEF) (Siegel et al., 2015) in monkeys performing an experimental task that required the animal to attend to either color or motion information while ignoring the non-cued feature of the stimuli. (c, d) RNNs with three different architectures, CTRNN, LowPassCTRNN, LSTM and three different nonlinearities, ReLU, ReTanh, Tanh are compared to neural datasets (see appendix B for details).
@@ 2 Related Work
Reviews. Recent reviews by Sucholutsky et al. (2023) and Klabunde et al. (2023) provide comprehensive overviews of representational similarity measures and their theoretical properties. While these reviews highlight the diversity of available metrics, they offer limited practical guidance on interpreting similarity scores.
Our work addresses this gap by proposing a general framework for evaluating similarity measures. By directly optimizing synthetic datasets to maximize their similarity to, for example, neural recordings, we can systematically investigate how different metrics prioritize various aspects of the data, such as specific principal components or task-relevant information. Similarity measures are often characterized by the invariance properties between representations with maximum similarity. In contrast, we focus on what drives intermediate similarity scores and how to interpret them, since we are often dealing with intermediate levels of similarity in practice when comparing models to brain data.
@@
However, our approach differs significantly in how we construct a diverse set of datasets with varying behavioral characteristics. Instead of relying on pre-trained models, we begin with unstructured noise and directly optimize it to maximize similarity to neural recordings. This allows us to address questions such as whether high similarity scores can be achieved without encoding task-relevant variables. Our optimization-based approach reveals how different similarity measures guide the emergence of task-relevant information within the synthetic datasets, offering a dynamic perspective on their properties. Additionally, by leveraging our optimization-based perspective for multiple similarity measures we can find the allowable ranges of scores, and identify dependencies between similarity measures.
@@ 3 Method
3.1 Measuring similarity
To evaluate the similarity of representations between two systems, we extract feature representations such as activity in a brain area or model layer in response to some sample stimuli. Our objective is to quantify the alignment between these representations using a similarity score. Assume two datasets $X$ and $Y$ represent these features with dimensions (sample, feature) and mean-centered columns. Datasets with temporal dynamics are reshaped from (time, sample, feature) to (time*sample, feature). We define a scoring function $\text{score}(X,Y)$ as a measure that increases with similarity, achieving a maximum of 1 when $X=Y$.
We consider the following similarity measures (see appendix C.1 for details):
@@ 3.1 Measuring similarity
  • Angular CKA is a variant of CKA that satisfies the axioms of a distance metric (Williams et al., 2021; Lange et al., 2022). It is defined as the arccosine of CKA. We additionally renormalize it to have a value between 0 and 1, where 1 is perfect similarity.
@@
    $\min_{Q\in\mathcal{O}}\text{arccos}\dfrac{\langle X,QY\rangle}{\|X\|_{F}\|Y\|_{F}}$
@@
  • Normalized Bures Similarity (NBS) (Tang et al., 2020) is the cosine of the angular Procrustes distance (Harvey et al., 2023). NBS is also closely related to CKA and mainly differs by a choice of matrix norm (see details in appendix C.1).
    $\text{NBS}(X,Y)=\dfrac{\|X^{T}Y\|_{*}}{\sqrt{\|XX^{T}\|_{*}\|YY^{T}\|_{*}}}$
  • Ridge Regression finds the best linear mapping $B$ that predicts a reference dataset $X$ from a dataset $Y$ (Yamins et al., 2014; Schrimpf et al., 2018). We measure the goodness of fit using $R^{2}$ and use ridge regularization as well as 5-fold cross-validation that tests generalization across different experimental conditions. We use $\lambda=100$ by default and show results with varying $\lambda$ in appendix A.
    $B^{*}=\operatorname*{arg\,min}_{B}\|X-YB\|^{2}_{F}+\lambda\|B\|_{F}$
    $R^{2}_{LR}=1-\dfrac{\|X-YB^{*}\|^{2}_{F}}{\|X\|^{2}_{F}}$
  • Linear Regression. Same as Ridge Regression but for $\lambda=0$ (unregularized).
3.2 Optimizing similarity scores
To better characterize similarity measures we optimize synthetic datasets $Y$ to become more similar to a reference dataset $X$. We initialize the synthetic dataset $Y$ by randomly sampling from a standard Gaussian distribution with the same shape as $X$. We use Adam (Kingma & Ba, 2017) to optimize $Y$ to maximize the similarity score with $X$, leveraging the differentiability of the similarity measures, and stop the optimization when the score reaches a fixed threshold near 1. Note that some similarity measures have parameters to optimize to compute the similarity score. Our method can be applied in such cases too, as long as the similarly score is differentiable with respect to the input datasets. For example, in the case of linear regression, we directly differentiate PyTorch’s lstsq function.
As we optimize $Y$ towards greater similarity with $X$, we simultaneously evaluate how well task-relevant variables can be linearly decoded from the evolving synthetic data. This decoding analysis employs logistic regression with stratified 5-fold cross-validation. For datasets with temporal dynamics, we fit a separate decoder for each time step and report the average accuracy across time.
We also evaluate how each principal component (PC) of the reference dataset $X$ is captured as $Y$ is optimized for similarity. Specifically, for each PC $v_{i}$ of $X$, we use linear regression to find the optimal linear combination of columns of $Y$ that best predicts $Xv_{i}$. The goodness of fit is then quantified using the $R^{2}$ coefficient:
$R_{PC_{i}}^{2}:=1-\dfrac{\|(X-\hat{X})v_{i}\|^{2}}{\|Xv_{i}\|^{2}}$
where $\hat{X}=Y\hat{B}$ and $\hat{B}=\text{argmin}_{B}\|X-YB\|^{2}_{F}$.
@@ 4 Results
4.1 What constitutes a good similarity score depends on the similarity measure and the dataset
What is a good value for a similarity score? Our approach to address this question is to examine, across five neural datasets, the similarity score required for a synthetic dataset to encode task relevant information to the same degree as the neural data. More specifically, the dots along the x-axis in Figure 3 indicate the score when a decoder trained on the synthetic data can extract 90% of the full task relevant information present in the neural data. Consider the Mante 2013 dataset, and notice that the “good” scores for the six similarity measures are quite different, ranging from less than 0.5 to almost 1. So even for the same dataset, the similarity score required to encode task relevant information to a similar extent as the neural data, can be very different across similarity measures.
Now if we look across the five neural datasets, at a single similarity measure, we can see that what constitutes a good score varies depending not only on the similarity measure but also on the dataset. An angular Procrustes score above 0.5 may constitute a good score for the Mante 2013 dataset but a score above 0.8 is required for the Siegel 2015 dataset (see also Supplementary Figure S1).
High similarity scores do not guarantee encoding of task relevant variables. A crucial point to make with Figure 3 is that high similarity scores near the maximum value of 1, particularly for CKA and unregularized linear regression without cross-validation, do not guarantee that models encode task-relevant information in a manner consistent with neural data, i.e. the CKA and linear regression curves in the Siegel 2015 dataset do not approach the horizontal line showing the decode accuracy for the neural data. There may be important features in a dataset that are not captured by a model even when the model-data similarity score is high.
For linear regression, even cross-validated and regularized, a high similarity score does not necessarily mean that the task relevant information is encoded in a similar manner to the neural data (Supplementary Figure S2).
Surprisingly, in Figure 3 some optimized datasets encode task relevant information better than the original neural data (colored lines showing decode accuracy are above the horizontal dashed line). This might suggest that the optimized datasets are denoised versions of the original data. We tested this hypothesis by removing low-variance principal components from the neural dataset, but this did not change the results. This suggests it is probably a more complex form of denoising and would require further investigation (Supplementary Figure S3).
Figure 3: What constitutes a good score varies depending on the similarity measure and the dataset. Decode accuracy for experimental variables versus similarity scores. The experimental variables are color vs motion contexts (binary variable) for Mante 2013 and Siegel 2015, reaching direction (total of 8 directions) for Hatsopoulos 2007, object categories (total of 8 categories) for MajajHong 2015, and texture vs noise categories (binary variable) for FreemanZiemba 2013. Horizontal dashed lines show the decode accuracy from the neural data (upper line) and chance level (lower line). Colored dots above the x-axis indicate the similarity scores when the decode accuracy reaches 90% midway between chance level and the decode accuracy from the reference neural dataset.
@@ 4.2.1 Low-dimensional synthetic datasets
Question. How much of the neural data must be captured by a synthetic dataset or model before the decode accuracy reaches the level seen in the neural dataset itself? One perspective on this question is to decode the task variables from neural data after projecting onto principal components 1 through N, where principal component 1 captures the most variance. As we might expect, in order to capture all the information about the task variables, at least several principal components must be included in the decode (Figures S1c and S1d show an example of this analysis for two of the neural datasets). This motivates the following hypothesis.
@@
Hypothesis. Perhaps the reason that some variations of linear regression and CKA similarity scores can be so high while the synthetic data fails to encode task variables is because these similarity measures preferentially rely on the top few principal components. In other words, if these measures only encode information from the top principal components then this may not be sufficient to encode all the task variables. We explore this hypothesis in the following set of analyses with a synthetic dataset based on the neural recordings from Mante et al. 2013.
@@
Figure 4: Different similarity measures differentially prioritize learning principal components of the data. (a) Reference dataset used as a target during optimization. (b, c) Initial Gaussian random noise data is updated to maximize similarity with the reference dataset, as quantified by one of the similarity measures. The transformation of the random noise dataset is shown at the top of panel b. The first principal component of the reference dataset is increasingly well captured by the optimized data as the similarity scores increase (yellow curves). The second, lower variance, component is also learned when maximizing the angular Procrustes similarity but is only captured at high similarity scores when maximizing linear regression, CKA, and angular CKA similarity. (d) Four reference datasets with decreasing variance along the second principal component. (e) Similarity measures capture both principal components when their variance is approximately equal. However, when the variance differs, CKA and linear regression preferentially neglect the low variance component (curves colored according to asymmetry of variance distribution).
Figure 4a shows the reference dataset (compare to Figure 2a). We can think of this reference dataset as a low-dimensional neural trajectory summarizing the population activity of many neurons, or alternatively, as the firing rates of two neurons over time (shown here encoding the two task variables of choice and dot motion coherence), recorded during six different experimental conditions, with the color in Figure 4a denoting the condition. Figure 4b shows the transformation of an initially random Gaussian noise dataset as it is optimized to maximize either the angular Procrustes or CKA similarity score with respect to the reference dataset. The score increases from an initial value near 0 to a maximum near 1 as optimization progresses, with the insets at the top of the figure showing the optimized noise dataset at various points during this procedure. The yellow curve shows how well the optimized dataset captures the first principal component of the reference dataset, as quantified by $R^{2}$, throughout optimization (see section 3.2 for details). Notice that the second principal component, shown in purple, is only captured at a much higher optimization score for CKA versus angular Procrustes. Figure 4c shows the same results when a synthetic Gaussian noise dataset is optimized towards the reference dataset using either linear regression similarity or angular CKA similarity (Williams et al., 2021).
@@
Dependence on the variance distribution. The optimization dynamics not only depend on the similarity measure but also on the variance distribution of the reference dataset. Figure 4d shows four reference datasets with the same variance along the first principal component but decreasing variance along the second. If both dimensions have approximately equal variance then angular Procrustes, CKA, and linear regression will learn both dimensions similarly during optimization as shown by the white curve in Figure 4e. The curves are colored to indicate the fraction of variance the second principal component has relative to the first, so 1 indicates both principal components have the same variance. As the asymmetry between the principal components grows, the optimization to maximize angular Procrustes similarity still effectively learns the lower variance principal component (first column, second row), while CKA and linear regression do not capture the second principal component of the reference dataset until much later during optimization.